evalute the function at the given value of the independent variable and simplify
g(x)=2x-4;g(x-1)

Answer by stanbon(57307)

g(x)=2x-4;g(x-1)
-----
g(x-1) = 2(x-1)-4 = 2x-6
==============================
cheers,
Stan H.
============